OpenStudy (anonymous):

Simplify (√12 -√2)(√3 +√18)

OpenStudy (anonymous):

yOU WOULD SIMPLIFY RADICAL 12 INTO RADICAL4 AND RADICAL 3 SIMPLIFY RADICAL 4 INTO 2 RADICAL 3 STAYS THE SAME THE NEW THING WOULD BE 2 RADICAL 3 rIGHT

OpenStudy (anonymous):

Yay your their

OpenStudy (anonymous):

*there

OpenStudy (anonymous):

√12√3+√12√18-√2√3-√2√18

OpenStudy (anonymous):

so you distribute right?

OpenStudy (anonymous):

sowhen I simplified I got \[6 + 6\sqrt{6} -\sqrt{6} -\sqrt{6}\] IS THAT RIGHT?

OpenStudy (raden):

for √2√18, it should be √36 = 6

OpenStudy (anonymous):

but there's a negative sign in front of the radical 2

OpenStudy (anonymous):

the last figure is wrong ....check it \[\sqrt{2}\sqrt{18}\]

OpenStudy (raden):

ye, just give sign - infront it, be -6 :)

OpenStudy (anonymous):

okay \[6 +6\sqrt{6} -\sqrt{6} -6\]

OpenStudy (anonymous):

Right?

OpenStudy (raden):

yes... now, cncel out for 6-6, be 0

OpenStudy (anonymous):

so its 5 radical 6 right ?

OpenStudy (raden):

agree

OpenStudy (anonymous):

Thanks!

OpenStudy (raden):

welcome